# Understanding Bias Reinforcement in LLM Agents Debate

### v4

@@ 5 DReaMAD: Diverse Reasoning via MAD
Average ReAct 0.87 0.36 0.21 0.22 -
+ CoT-Prompting 0.90 0.40 0.30 0.20 -
@@
5.1 Strategic Prior Knowledge Elicitation
To address strongly consistent bias, DReaMAD integrates a structured module that ensures systematic extraction and refinement of the LLM’s internal strategic knowledge before the debate. First, the model is prompted to reinterpret the given problem, leading to a more organized understanding of the strategic context (Game Situation Reinterpretation). Next, it formulates a set of high-level strategies that can be applied to the scenario at hand (General Strategy Formulation), preventing the model from settling too early on potentially flawed reasoning. As shown in Table 3, this module enhances the model’s performance in tasks that require strategic prior knowledge (Figure 4-1). We set the temperature hyperparameter to be 0.1 for strategic consistency.
5.2 Perspective Diversification
Building on MAD, DReaMAD mitigates argument homogenization by ensuring each agent adopts a distinct viewpoint prior to the debate. This approach is inspired by the Learning from Multiple Approaches concept in education theory (Council et al., 2005; Cleaves, 2008), which suggests that individuals improve their problem-solving skills by exploring multiple representations of the same problem. Analogously, this module ensures that each agent is given differentiated initial prompts that guide reasoning along distinct strategic trajectories. By self-customizing initial prompts for each model instance, DReaMAD encourages unique strategic perspectives and reduces the risk of bias reinforcement (Figure 4-2) as demonstrated in Figure 5 and Figure 6 (left). In Figure 5, although each agent receives the same initial prompt, they independently generate different optimal prompts, leading to distinct distributions of reasoning. This process reduces each agent’s bias and fosters a more robust debate. Remarkably, even in a state with very strong consistency, the discussion converges well toward the correct reasoning direction. Here, we set temperature to be 0.7 for diversity.
@@ 6 Experiments
Table 4: Winning rate comparison across different models and different self-correction methods. This is an result based on MetaNIM Arena simulator. The best results are highlighted in bold.
LLM Models Prompting Methods NIM Fibonacci Kayles Chomp C.Queen
Normal Misère Normal Misère Single 2 Rows Rectangular Square Normal
GEMINI-1.5-flash
Standard Prompting 0.32 0.54 0.16 0.10 0.54 0.56 0.78 0.18 0.46
+ ReAct 0.10 0.68 0.16 0.76 0.50 0.30 0.42 0.12 0.46
+ Self-Refinement 0.14 0.66 0.18 0.36 0.50 0.46 0.46 0.16 0.42
+ Self-Consistency 0.04 0.28 0.28 0.86 0.30 0.24 0.74 0.0 0.34
+ MAD 0.06 0.30 0.12 0.78 0.54 0.20 0.74 0.14 0.58
+ MAD2 0.26 0.26 0.08 0.06 0.44 0.36 0.68 0.10 0.58
+ DReaMAD 0.38 0.84 0.16 0.94 0.58 0.62 0.60 0.22 0.74
GPT-4o-mini Standard Prompting 0.38 0.54 0.22 0.28 0.46 0.48 0.46 0.38 0.12
+ ReAct 0.22 0.68 0.20 0.34 0.40 0.26 0.58 0.34 0.28
+ Self-Refinement 0.22 0.70 0.18 0.50 0.46 0.52 0.52 0.44 0.24
+ Self-Consistency 0.14 0.52 0.34 0.46 0.32 0.20 0.54 0.26 0.30
+ MAD 0.28 0.62 0.22 0.64 0.42 0.28 0.52 0.56 0.44
+ MAD2 0.34 0.20 0.18 0.18 0.50 0.34 0.44 0.90 0.14
+ DReaMAD 0.98 0.74 0.54 0.72 0.68 0.84 0.64 0.22 0.76
@@
We utilized the benchmark MetaNIM Arena as both our dataset and simulator, as it provides a controlled environment for evaluating reasoning under grounded strategic tasks. Our investigation focuses on three key questions: (1) Does our approach improve reasoning quality compared to existing prompting techniques? (2) Does our approach prove its strategic reasoning quality under adversarial decision-making environments? (3) Does generating diverse prompts contribute to better decision-making within the debate framework?
@@
To address these questions, we compare DReaMAD with standard prompting methods including ReAct (Yao et al., 2023), Chain-of-Thought (CoT), Self-Consistency (Wang et al., 2022), Self-Refinement (Madaan et al., 2024), and Multi-Agent Debate (MAD (Du et al., 2023), MAD2 (Liang et al., 2023)). The details of standard and CoT prompts are provided in Appendix B.
@@
Our method builds on the MAD framework by Du et al. (2023), augmenting it with structured self-prompt refinement and perspective diversification. For self-refinement, we follow the methodology of Madaan et al. (2024), applying three iterative refinement steps. Similarly, for MAD, we conducted up to three rounds of debate, following Du et al. (2023), with the process terminating early if a consensus is reached before the final round.
We also investigate whether the observed improvements generalize across different LLM architectures, including both GPT and GEMINI models.
Table 5: Accuracy (%) on math-reasoning benchmarks. Columns group the five algorithms for each backbone model. Bold indicates the best algorithm for a given (dataset, model) pair.
@@
ReAct Self-Refinement Self-Consistency MAD DReaMAD ReAct Self-Refinement Self-Consistency MAD DReaMAD
AIME 2024 76.7 73.3 86.7 73.3 90.0 0.0 3.3 10.0 3.3 10.0
AMC 2023 97.5 100 100 100 100 60.0 52.5 52.5 60.0 62.5
Table 6: Accuracy (%) on CommonsenseQA dataset. Bold indicates the best performance. We abbreviate Self-Refinement as Self-Refine. and Self-Consistency as Self-Consist.
@@
GPT-4o 83.6 49.2 83.6 82.0 84.4
GPT-4o-mini 78.7 61.5 78.7 75.4 79.5
6.1 Does DReaMAD Improve Reasoning Quality?
This experiment isolates the effect of strategic prior knowledge elicitation (§5.1), allowing us to assess whether our method enhances decision-making without relying on debate dynamics.
Setup. To evaluate the effectiveness of our approach in improving reasoning capabilities, we compare DReaMAD without the debate process against ReAct and zero-shot CoT prompting across multiple models in the MetaNIM Arena dataset (§A.3). For showing versatility of DReaMAD, we conduct experiments on four variants of LLMs as shown in Table 3.
Results. Table 3 demonstrates that DReaMAD consistently outperforms both ReAct and CoT prompting across all models and tasks. These results highlight the impact of our method in reinforcing structured strategic reasoning, even without the iterative correction process of debate. Notably, our approach leads to substantial improvements in NIM, Fibonacci, and Kayles, which are environments where long-term strategic planning plays a crucial role. Since defining a general winning strategy in Chomp is non-trivial, applying prior knowledge is challenging and results in less effectiveness compared to other games. Furthermore, we observe that models with inherently weaker reasoning abilities benefit the most from strategic prior knowledge elicitation (e.g., GPT-4o-mini with +17%p, GEMINI-1.5-flash with +12%p on average).
6.2 DReaMAD in Adversarial Strategic Decision-Making
Setup. We applied DReaMAD to GPT-4o-mini and GEMINI-1.5-flash and compared it with self-correction methods, including standard-prompt, ReAct, self-refinement, self-consistency, and MAD. To demonstrate its effectiveness, we used GPT-4o as the opponent model due to its superior performance. In this experiments, we utilized MetaNIM Arena simulator (§A.4) to maximize the effect of generating diverse prompts. We aimed to validate our hypothesis in a simulator that requires strategic decision-making within complex dynamics. We ran 50 independent episodes and average the win-rate.
Results. As shown in Table 4, DReaMAD consistently outperforms other self-correction methods across various strategic environments, demonstrating a significant improvement in winning rates. This result suggests that our approach enables LLM agents to effectively adapt to complex dynamics, particularly in adversarial decision-making scenarios where strategic reasoning is crucial. However, we observe that DReaMAD struggles in the Chomp game, which aligns with our hypothesis that Chomp lacks a well-defined generalized winning strategy. Unlike other tested environments, Chomp requires more exploratory play rather than direct reasoning from prior knowledge, highlighting a limitation of our method in environments where strategic heuristics are less structured.
6.3 Does Generating Diverse Prompts Improve Performance?
We assess whether prompt diversity improves decision quality within the MAD framework. To this end, we compare two settings: (1) identical prompts generated via the Strategic Prior Knowledge Elicitation module for both agents ((+)MAD), and (2) distinct, self-generated prompts per agent, as in DReaMAD (Figure 6, left). Experiments were conducted on four variants of the MetaNIM Arena simulator, NIM (Normal and Misère) and Fibonacci (Normal and Misère), over 50 episodes each (§A.4). Our results show that incorporating diverse prompts within the MAD framework significantly enhances performance, validating the effectiveness of our Perspective Diversification module.
We also examine the effect of sampling temperature on prompt diversity in the Fibonacci task. Within both the Strategic Prior Knowledge Elicitation and Perspective Diversification modules, we vary the temperature from 0.0 to 1.0. As shown in Figure 6 right, higher temperature (further diversity) correlates with increased optimal action accuracy, indicating that greater diversity in generated prompts contributes to improved reasoning performance.
6.4 Generalization to Math and Commonsense Reasoning
While our primary benchmark focuses on structured games, we further evaluate DReaMAD on NLP tasks to test its broader applicability. Specifically, we consider algebra and number theory problems from AIME 2024 and AMC 2023, as well as CommonsenseQA (Talmor et al., 2018), which require symbolic reasoning and multi-step inference. These tasks are representative of domains where chain-of-thought reasoning is essential, making them suitable for evaluating generalization. We also examine whether DReaMAD benefits reasoning-specialized models, such as GPT-o3-mini, in a similar manner as general-purpose LMs. Experiments are conducted using standard accuracy metrics across model-task pairs. Results in Table 5 and Table 6 show that DReaMAD not only boosts accuracy across heterogeneous tasks but also yields clear gains on reasoning-oriented models such as GPT-o3-mini, underscoring the method’s robustness well beyond the structured-game domain.
7 Conclusions

### v5

@@ 4 DReaMAD: Counteracting Belief Entrenchment
Qwen3-4B ReAct 0.99 $\pm$ 0.02 0.42 $\pm$ 0.09 0.08 $\pm$ 0.05 0.47 $\pm$ 0.03 0.49
+ CoT-Prompting 0.98 $\pm$ 0.02 0.38 $\pm$ 0.04 0.08 $\pm$ 0.03 0.48 $\pm$ 0.02 0.48
@@
Average ReAct 0.89 0.37 0.18 0.27 -
+ CoT-Prompting 0.92 0.39 0.26 0.26 -
[…]
To address the limitations of standard MAD identified in Section 3, we introduce DReaMAD (Diverse Reasoning via Multi-Agent Debate with Refined Prompt). Our framework counteracts belief entrenchment by targeting its two root causes: the initial static bias ($p^{(0)}$) and the homogenized debate dynamics.
4.1 Stage 1: Strategic Prior Knowledge Elicitation (SPKE)
The first stage aims to correct the model’s initial reasoning bias before any debate begins. Standard prompting often leads models to rely on superficial heuristics, resulting in a skewed initial probability distribution $p^{(0)}$.
Method.
We implement a two-step elicitation process. First, in Game Situation Reinterpretation, the model is prompted to analyze the problem’s structure and constraints explicitly. Second, in General Strategy Formulation, it formulates high-level winning strategies. This step stops the model from locking into a wrong path early. We set the temperature to 0.1 to ensure consistent strategic output.
Theoretical Justification.
By forcing the model to ground its reasoning in domain-specific strategic knowledge, SPKE shifts the initial probability mass $p^{(0)}$ towards the optimal action. According to our model in Eq. 1, increasing the initial probability of the correct argument is crucial because the subsequent debate dynamics tend to amplify the dominant initial belief.
4.2 Stage 2: Perspective Diversification
Even with a corrected initial belief, standard MAD suffers from an “echo chamber” effect where homogeneous agents reinforce each other’s remaining biases. This corresponds to the persuasiveness being dependent on the argument’s initial popularity rather than its quality.
Method.
We break this dependency by assigning each agent a unique, self-generated prompt. Inspired by the Learning from Multiple Approaches theory (Council et al., 2005; Cleaves, 2008), this encourages agents to adopt distinct reasoning paths. For example, as shown in Figure 5, even starting from the same state, agents guided by diverse prompts generate distinct internal reasoning, preventing premature convergence to a wrong consensus. We set the temperature to 0.7 to promote diversity.
Theoretical Justification.
This module directly targets the debate dynamics. By forcing agents to adopt different perspectives, we ensure arguments win based on logic, not popularity $p^{(0)}$. This ensures that arguments are evaluated based on their intrinsic logical merit rather than their alignment with the majority bias, fostering a robust, truth-seeking debate.
After these stages, the agents conduct a structured multi-agent debate. The complete workflow is illustrated in Figure 1 and the detailed prompt formulation used in these two modules is documented in the Appendix A.2.2.
@@ 5 Experiments
Figure 6: Winning rate comparison across different models and different self-correction methods. This is a result based on MetaNIM Arena simulator. The best results are highlighted in bold. We report win rates with 95% confidence intervals over 50 episodes.
@@
We utilized the benchmark MetaNIM Arena as both our dataset and simulator, as it provides a controlled environment for evaluating reasoning under grounded strategic tasks. Our investigation focuses on three key questions: (1) Does our approach improve reasoning quality compared to existing prompting techniques? (2) Does our approach prove its strategic reasoning quality under adversarial decision-making environments? (3) Does generating diverse prompts contribute to better decision-making within the debate framework? Further, we conduct two more ablation studies: (4) Can DReaMAD achieve long CoT reasoning without further training? (5) Can DReaMAD be generalized to general NLP domain?
@@
To address these questions, we compare DReaMAD with standard prompting methods including ReAct (Yao et al., 2023), Zero-shot Chain-of-Thought (CoT), Self-Consistency (Wang et al., 2022), Self-Refinement (Madaan et al., 2024), and Multi-Agent Debate (MAD (Du et al., 2023) and Multi-Persona Debate (MPD) (Liang et al., 2023)). The details of standard and CoT prompts are provided in Appendix A.2.2.
@@
Our method builds on the MAD framework by Du et al. (2023), augmenting it with structured self-prompt refinement and perspective diversification. For self-refinement, we follow the methodology of Madaan et al. (2024), applying three iterative refinement steps. Similarly, for MAD, we conducted up to three rounds of debate, following Du et al. (2023), with the process terminating early if a consensus is reached before the final round. We also investigate whether the observed improvements generalize across different LLM architectures, including GPT Achiam et al. (2023), Gemini Team et al. (2023), and Qwen Yang et al. (2025) models.
5.1 Does DReaMAD Improve Reasoning Quality?
This experiment isolates the effect of strategic prior knowledge elicitation, allowing us to assess whether our method enhances decision-making without relying on debate dynamics.
Setup. To evaluate the effectiveness of our approach in improving reasoning capabilities, we compare DReaMAD without the debate process against ReAct and Zero-shot CoT prompting across multiple models in the MetaNIM Arena dataset (§A.2). For showing versatility of DReaMAD, we conduct experiments on five variants of LLMs as shown in Table 1.
[…]
5.2 DReaMAD in Adversarial Strategic Decision-Making
Setup. We applied DReaMAD to GPT-4o-mini, Gemini-1.5-flash, and Qwen3-4B then compared it with self-correction methods, including standard-prompt, ReAct, self-refinement, self-consistency, MAD and MPD. To demonstrate its effectiveness, we used GPT-4o as the opponent model due to its superior performance. In these experiments, we utilized MetaNIM Arena simulator to maximize the effect of generating diverse prompts. We aimed to validate our hypothesis in a simulator that requires strategic decision-making within complex dynamics. We ran 50 independent episodes and average the win-rate.
Results. As shown in Figure 6, DReaMAD consistently outperforms other self-correction methods across various strategic environments, demonstrating a significant improvement in winning rates. This result suggests that our approach enables LLM agents to effectively adapt to complex dynamics, particularly in adversarial decision-making scenarios where strategic reasoning is crucial. Notably, even for a small model such as Qwen3-4B, DReaMAD achieves the best overall performance among the compared methods; however, the magnitude of improvement tends to be larger for higher-capacity models, indicating that additional model capacity better translates elicited strategic priors into effective decision policies. However, we observe that DReaMAD struggles in the Chomp game, which aligns with the fact that, although a winning strategy is guaranteed to exist for the first player by the strategy-stealing argument, no general closed-form solution is known for arbitrary board sizes. This proof establishes the existence of a winning strategy without explicitly constructing it, making the task highly non-trivial in practice. The absence of a fully characterized optimal strategy means that effective play often requires exploratory search rather than direct reasoning from prior knowledge. This highlights a limitation of our method in environments where strategic heuristics are only partially understood or insufficiently structured.
5.3 Does Generating Diverse Prompts Improve Performance?
We assess whether prompt diversity improves decision quality within the MAD framework. To this end, we compare two settings: (1) identical prompts generated via the Strategic Prior Knowledge Elicitation module for both agents as (+)MAD, and (2) distinct, self-generated prompts per agent, as in DReaMAD (Figure 7, left). Experiments were conducted on four variants of the MetaNIM Arena simulator, NIM (Normal and Misère) and Fibonacci (Normal and Misère), over 50 episodes each. Our results show that incorporating diverse prompts within the MAD framework significantly enhances performance, validating the effectiveness of our Perspective Diversification module.
We also examine the effect of sampling temperature on prompt diversity in the Fibonacci task. Within both the Strategic Prior Knowledge Elicitation and Perspective Diversification modules, we vary the temperature from 0.0 to 1.0. As shown in Figure 7 right, higher temperature (further diversity) correlates with increased optimal action accuracy, indicating that greater diversity in generated prompts contributes to improved reasoning performance. However, we also note that excessively high temperatures could risk degrading reasoning coherence by introducing noise. Determining the optimal temperature that balances diversity and logical consistency remains an important consideration for practical applications.
@@ 5.3 Does Generating Diverse Prompts Improve Performance?
ReAct Self-Refine. Self-Consist. MAD DReaMAD ReAct Self-Refine. Self-Consist. MAD DReaMAD
AIME 2024 72.0 $\pm$ 4.47 74.0 $\pm$ 1.49 82.7 $\pm$ 2.79 82.7 $\pm$ 2.79 82.7 $\pm$ 2.79 8.67 $\pm$ 4.47 8.67 $\pm$ 2.98 10.7 $\pm$ 2.71 11.3 $\pm$ 5.06 12.0 $\pm$ 5.58
AMC 2023 99.0 $\pm$ 1.37 98.0 $\pm$ 1.12 99.5 $\pm$ 1.12 99.5 $\pm$ 1.12 100 $\pm$ 0.0 51.5 $\pm$ 3.29 54.5 $\pm$ 4.81 58.0 $\pm$ 5.42 55.0 $\pm$ 5.59 59.5 $\pm$ 4.47
Table 2: Accuracy (%) on math-reasoning benchmarks. We report the mean and standard deviation over 5 independent runs. Bold indicates the best algorithm for a given pair. While DReaMAD consistently achieves the highest mean accuracy, the overlap in standard deviations with Self-Consistency suggests comparable performance in some settings, highlighting the robustness of ensemble-based methods.
Figure 8: Performance gain of DReaMAD when applied to the Non-thinking (N/T) mode. The figure shows the average win rate over 50 runs for the Qwen3-1.7B (top) and Qwen3-4B (bottom) models. The notation NIM (n, k) indicates a game starting with n stones and a maximum take of k stones per turn.
5.4 Achieving Long CoT Reasoning Without Specialized Training
Researchers often assume that long CoT reasoning requires costly post-training, such as the multi-stage reinforcement learning pipeline used to create Qwen3’s thinking mode. In this section, we investigate a critical question: can the latent reasoning capabilities of a base model (non-thinking mode) be unlocked to match this long CoT reasoning performance without undergoing such training?
To explore this, we analyze the performance of the Qwen3-1.7B and 4B model (Yang et al., 2025) on 4 variants of NIM-Normal game which have different initial number of stones and maximum number of stones that can be taken under three conditions, as shown in Figure 8: (1) the standard, Non-thinking (N/T) mode; (2) the heavily post-trained Thinking mode; and (3) DReaMAD on top of the N/T mode.
The results reveal the substantial value gained by our approach, demonstrating its effectiveness across different model scales. The standard Non-thinking mode models establish a performance baseline, which our framework significantly improves upon. For the Qwen3-1.7B model, DReaMAD achieves a remarkable 226% average performance gain, elevating a weak baseline to 81.3% of the specialized Thinking mode’s capability.
This effect is even more pronounced on the more capable Qwen3-4B model. Here, applying DReaMAD yields a 45.1% average performance gain, bringing the base model to 97.9% of the Thinking mode’s performance and thus achieving near-parity: while the Thinking mode still attains the highest raw accuracy, DReaMAD recovers most of its benefit without any additional post-training. The primary gain is one of efficiency and accessibility; DReaMAD provides a training-free pathway to elicit long CoT reasoning from general-purpose models at test-time, bypassing the need for separate, resource-intensive training pipelines.
5.5 Generalization to Math and Common Sense Reasoning
@@
GPT-4o 83.93 $\pm$ 1.24 83.11 $\pm$ 1.37 84.75 $\pm$ 0.93 84.26 $\pm$ 1.47 83.93 $\pm$ 1.10
GPT-4o-mini 77.38 $\pm$ 1.37 75.25 $\pm$ 3.95 76.89 $\pm$ 0.69 77.38 $\pm$ 0.73 78.85 $\pm$ 1.07
Table 3: Accuracy (%) on CommonsenseQA dataset. We report the mean and standard deviation over 5 runs. Bold indicates the best performance. DReaMAD shows a statistically significant improvement over ReAct for GPT-4o-mini, while performing comparably to Self-Consistency for GPT-4o.
While our primary benchmark focuses on structured games, we further evaluate DReaMAD on NLP tasks to test its broader applicability. Specifically, we consider algebra and number theory problems from AIME 2024 and AMC 2023, as well as CommonsenseQA (Talmor et al., 2018). Experiments are conducted using standard accuracy metrics across model-task pairs, averaged over 5 independent runs.
Results in Table 2 and Table 3 show that DReaMAD remains competitive or superior beyond our MetaNIM Arena benchmark. In math reasoning (Table 2), DReaMAD clearly improves over ReAct and Self-Refine. For example, on AIME 2024 with GPT-o3-mini it reaches 82.7%, about 9–11 points higher than ReAct (72.0%) and Self-Refine (74.0%), and on AMC 2023 it attains 100% accuracy, slightly surpassing Self-Consistency. For GPT-4o, DReaMAD also gives the highest means on both AIME 2024 and AMC 2023, though the confidence intervals overlap with Self-Consistency and MAD, indicating comparable robustness.
Similarly, in CommonsenseQA (Table 3), DReaMAD achieves the highest mean accuracy for GPT-4o-mini (78.85%), demonstrating statistically significant gains over baselines. For the stronger GPT-4o model, it performs on par with Self-Consistency. These findings align with our observations in Section 5.4 (Qwen analysis): while specialized training (e.g., Thinking mode) sets a high baseline, DReaMAD offers a training-free pathway to unlock latent reasoning capabilities, achieving long CoT performance or further refining the outputs of reasoning-specialized models.
6 Related Works
6.1 Cognitive Bias and Belief Entrenchment in LLMs
While the term bias in Large Language Models (LLMs) frequently refers to societal or demographic disparities (Gallegos et al., 2024), our work investigates cognitive biases, which are systematic deviations from rational decision-making that hinder robust reasoning. Specifically, we focus on belief entrenchment, a phenomenon where models exhibit excessive confidence in their initial, often erroneous, outputs and resist correction.
Recent studies suggest that LLMs often prioritize consistency over accuracy. This tendency arises because models are trained to maximize the likelihood of the next token based on context, leading them to align responses with user prompts or their own previous outputs rather than objective truth (Perez et al., 2023). Furthermore, Turpin et al. (2023) demonstrated that Chain-of-Thought (CoT) reasoning is frequently biased by the model’s initial prediction. Instead of reasoning to find the answer, the model generates explanations to justify its initial choice. This aligns with our finding that models tend to reinforce their static belief, defined as the initial probability distribution over actions, regardless of its correctness.
This belief entrenchment is further intensified by the model’s reliance on consistency as a proxy for correctness. While Wang et al. (2022) showed that aggregating consistent reasoning paths generally improves performance, this approach fails when the model is confident in a wrong answer, as shown. In such cases, the model consistently generates plausible but incorrect responses. Although models possess some capability to evaluate the validity of their own claims (Kadavath et al., 2022), this internal calibration often fails to override the reinforced consistency in iterative settings. Consequently, methods like Self-Refinement often fail to fix errors (Huang et al., 2024; Valmeekam et al., 2022), instead creating “reasoning loops” that amplify the model’s inherent biases. Our work extends these observations to the multi-agent setting, demonstrating that without structural intervention to diversify perspectives, Multi-Agent Debate (MAD) acts as a dynamic amplifier of these static cognitive biases, effectively forming an echo chamber of homogeneous agents.
6.2 Prompt Engineering and Self-Correction in LLMs
Prompt engineering shapes model outputs without retraining, potentially improving generalization and reducing bias (Brown et al., 2020; Reynolds and McDonell, 2021; Zhao et al., 2024; Schulhoff et al., 2024; Shin et al., 2024). However, fully eliminating biases in complex reasoning remains challenging (Jiang et al., 2022; Lu et al., 2022).
Meanwhile, self-correction mechanisms in LLMs refine responses without external supervision (Ganguli et al., 2023; Liu et al., 2024; Kamoi et al., 2024). Self-consistency, for instance, ensembles multiple outputs but converges on frequent rather than correct answers (Wang et al., 2022; Chen et al., 2023), and self-refinement can reinforce rather than fix biases (Wan and others, 2023; Shinn et al., 2023; Madaan et al., 2024; Huang et al., 2024). Feedback-loop methods such as STaR (Zelikman et al., 2022), Reflexion (Shinn et al., 2023), and SCoRe (Kumar et al., 2024) also struggle to reliably correct biases or foster diverse reasoning (Guo et al., 2024).
6.3 Multi-Agent Debate in LLMs
Multi-Agent Debate (MAD) enables LLM agents to critique each other, enhancing reasoning on complex tasks (Liang et al., 2023; Du et al., 2023). ChatEval (Chan et al., 2023), a multi-agent evaluation system, simulates human judgment to assess model output quality. Optimizations include task-specific strategies for improving debate effectiveness (Smit et al., 2024) and ACC-Debate, an actor-critic framework that trains models to specialize in debates, achieving benchmark gains (Estornell et al., 2024). Highlighting similar risks, Estornell and Liu (2024) provided a theoretical framework demonstrating that multi-LLM debates are susceptible to “echo chamber" effects, where a flawed majority opinion arising from shared misconceptions can suppress correct minority viewpoints. While these enhancements improve performance, studies reveal a key limitation: static evaluations focus on assessing predefined problems, whereas real-world decision-making often involves dynamic, interactive environments where biases can evolve. Understanding how biases shift in these settings is crucial for developing robust strategies that extend beyond conventional static benchmarks.
7 Discussion
Limitation. While our approach demonstrates consistent improvements across a variety of strategic reasoning tasks, it is not without limitations. For instance, in the game of Chomp, a winning strategy is guaranteed for the first player by the strategy-stealing argument, but no general, closed-form solution is known. Our framework, which relies on eliciting and refining prior strategic knowledge, is consequently less effective in such scenarios. This was reflected in our experiments, where performance on the Chomp dataset and simulator was significantly lower than on other games (Table 1 and Figure 6). Chomp aside, we also observe that while our method improves performance on CommonsenseQA, the gains are relatively modest compared to those on strategic and mathematical tasks (Table 2). This highlights that our method excels when a strong strategic foundation can be leveraged.
Future Directions. A key future direction involves comparing our single-model approach with heterogeneous model ensembles. While employing different models could inherently increase perspective diversity, this often introduces significant practical overhead. Our work prioritizes a more tractable solution by eliciting diversity from a single model, a crucial step for real-world deployment. Systematically evaluating the trade-offs between these two strategies, and potentially integrating them, remains a promising area for future research.
8 Conclusion